OpenStudy (anonymous):

Perform the indicated operation g(x) = x3 - 2 h(x) = 2x - 1

OpenStudy (igreen):

You can graph at https://www.desmos.com/calculator

OpenStudy (igreen):

Graph both of them, look for the point of intersection and that will be your solution.

OpenStudy (muzzack):

Question is \[g(x) = x3 - 2 \] is \[g(x) = x^3 - 2 \]

OpenStudy (igreen):

Yes, of course.. @Muzzack

OpenStudy (anonymous):

im confused i never did this before